OpenStudy (et365):

So, how do I work the integral(0,6) |3x-6|dx ?

OpenStudy (anonymous):

\[\int\limits_{0}^{6}(3x-6)dx\] is this ur integral?

OpenStudy (anonymous):

if so, you can start integrating. \[\int\limits_{0}^{6}3xdx-\int\limits_{0}^{6}6dx\]

OpenStudy (freckles):

\[|3x-6|=3x-6 \text{ if } 3x-6 \ge 0 \\ |3x-6|=-(3x-6) \text{ if } 3x-6<0 \\ \text{ solving the inequality part } \\ |3x-6|=3x-6 \text{ if } x \ge 2 \\ |3x-6|=-(3x-6) \text{ if } x<2 \\ \text{ so you have } \int\limits_0^6 |3x-6| dx=\int\limits_0^2 -(3x-6) dx+\int\limits_2^6 (3x-6) dx\]

OpenStudy (et365):

Yeah thats it freckles, the integral was the absolute value. Thanks guys!